# A phase transition for the two-dimensional random field Ising/FK-Ising model

### v1

@@ 4 Absolute continuity in weak-disorder regime
is close to 1 for typical disorder where
@@
In Section 4.1, we show the concentration of $Z(h)$ using a similar technique as in [8, Theorem 3.14] and in Section 4.2, we apply Lemma 3.1 to show the absolute continuity.
@@ 4.1 Concentration of $Z(h)$
Let $\Omega$ be a region such that $\Lambda_{N}\subset\Omega\subset\Lambda_{2N}$ and $\gamma\in\{-1,1\}^{\partial_{int}\Omega}$ be a boundary condition on $\partial_{int}\Omega$. Let $\mathcal{H}^{0}\subset{\mathbb{R}}^{\Omega}$ denote the collection of the external field $h$ such that
$\Bigg{|}\Bigg{\langle}~{}\prod_{v\in\Omega}\Bigg{[}1+\sigma_{v}\tanh\left(\frac{\epsilon h_{v}}{T}\right)\Bigg{]}~{}\Bigg{\rangle}^{\gamma,0}_{T,\Omega}-1\Bigg{|}\leq\sqrt{\epsilon N^{\beta(T)}}.$ (4.4)
Then $\mathcal{H}^{0}$ is measurable under the probability measure $\mathbb{P}$.
@@
Of course, the set $\mathcal{H}^{0}$ relies on $\gamma,\Omega,T$ and $\epsilon,$ but we omit them from the notation for concision. The same conventions will be used several times without further notice.
@@
Take $\Omega=\Lambda_{N},$ then there exist constants $c_{1},c_{2}>0$ depending only on $T$ such that $\mathbb{P}(\mathcal{H}^{0})\geq 1-c_{1}\exp\left(-c_{1}^{-1}\sqrt{\epsilon^{-1}N^{-\beta(T)}}\right)$ for any disorder strength $\epsilon\leq c_{2}N^{-\beta(T)}$ and any boundary condition $\gamma$ on $\partial_{int}\Lambda_{N}$.
@@
The case that $T=T_{c}$ is just [11, Lemma 3.4]. The case that $T<T_{c}$ and $T>T_{c}$ are similar. We just point out the difference here. Note that
@@
To control the right-hand side, using [11, Lemma B.1], the core is to bound
@@
For $T>T_{c}$, by the Edward-Sokal coupling between the Ising model and the FK-Ising model, we have
$\langle\sigma^{J}\rangle^{\gamma,0}_{T,\Lambda_{N}}=\phi^{\gamma,0}_{p,\Lambda_{N}}(\mathtt{Even}_{J}),$
where $\mathtt{Even}_{J}$ stands for the event that each cluster in $\omega^{\gamma}$ contains an even number of points in $J$.
@@
For any $I\subset\Lambda_{N}$, let $d_{x}=\frac{1}{2}\cdot dist(x,\partial_{int}\Lambda_{N}\cup I\setminus\{x\})$ and let $\mathcal{R}_{x}=\Lambda_{d_{x}}(x)$. Then it is easy to get that
$\mathtt{Even}_{J}\subset\bigcap_{x\in I}\{x\longleftrightarrow\partial_{int}\mathcal{R}_{x}\}.$ (4.6)
@@
$\langle\sigma^{J}\rangle^{\gamma,0}_{T,\Lambda_{N}}\leq\prod_{x\in I}\phi_{p,\mathcal{R}_{x}}^{\mathrm{w},0}\left(\{x\longleftrightarrow\partial_{int}\mathcal{R}_{x}\}\right)\leq\exp{\Big{(}-\sum_{x\in I}c\cdot d_{x}\Big{)}}.$ (4.7)
Then we get from the calculation in [8, Lemmas 8.1 and 8.3] that
@@
If $T\geq T_{c}$ and $\gamma=\mathrm{f}$ is the free boundary, then we get from CBC that
$\langle\sigma^{J}\rangle^{\mathrm{f},0}_{T,\Omega}\leq\langle\sigma^{J}\rangle^{\mathrm{w},0}_{T,\Lambda_{2N}}.$
The desired bound on $A_{k}$ then follows from the proof of Lemma 4.3. ∎
@@
Let $\mathcal{H}^{1}$ denote the set of external field that the following (P1) property holds:
@@
There exists a constant $c_{1}>0$ depending only on $T$ such that for any disorder strength $\epsilon\leq c_{1}N^{-\frac{1}{2}}$, we have $\mathbb{P}(\mathcal{H}^{1})\geq 1-c_{2}\exp\left(-c_{2}^{-1}\epsilon^{-1}N^{-\frac{1}{2}}\right)$.
@@
Recall that $f(x):=\log\cosh(x)$. Given the inequality $\frac{x^{2}-x^{4}}{2}\leq f(x)\leq\frac{x^{2}}{2}$, we calculate that
@@
Combining (4.3), (4.4) and (4.9), we obtain that for $h\in\mathcal{H}^{0}\cap\mathcal{H}^{1},$
$\left(1-\sqrt{\epsilon N^{\beta(T)}}\right)\left(1-\sqrt{\epsilon^{2}N}\right)$ $\leq Z(h)\cdot\exp\left(\frac{\epsilon^{2}N^{2}}{2T^{2}}\right)$ $\leq\left(1+\sqrt{\epsilon N^{\beta(T)}}\right)\left(1+\sqrt{\epsilon^{2}N}\right).$ (4.11)
@@ 4.2 Concentration of $F(h,\omega)$
Therefore, we can consider the “typical” external field such that $F(h,\omega)$ concentrates well.
@@
We denote by $\mathcal{H}^{2}$ and $\mathcal{H}^{3}$ the sets of the external field $\{h_{v}:v\in\Lambda_{N}\}$ that the following properties (P2) and (P3) hold respectively.
$(\text{\emph{P}}2):\quad\phi_{p,\Lambda_{N}}^{\gamma,0}\left(F(h,\omega)-\frac{2\epsilon^{2}N^{2}}{T^{2}}\leq\epsilon N^{\alpha(T)}\right)\geq 1-\exp\left(-\epsilon^{-\frac{1}{4}}N^{-\frac{\alpha(T)}{4}}\right)\,,$ (4.13)
$(\text{\emph{P}}3):\quad\phi_{p,\Lambda_{N}}^{\gamma,0}\left(F(h,\omega)-\frac{2\epsilon^{2}N^{2}}{T^{2}}\geq-\epsilon N^{\alpha(T)}\right)\geq 1-\exp\left(-\epsilon^{-\frac{1}{4}}N^{-\frac{\alpha(T)}{4}}\right)\,.$ (4.14)
@@
$\mathbb{P}(\mathcal{H}^{2}\cap\mathcal{H}^{3})\geq 1-c_{2}\exp\big{(}-c_{2}^{-1}\epsilon^{-\frac{1}{2}}N^{-\frac{\alpha(T)}{2}}\big{)}.$
@@
The proof is essentially an application of the Markov’s inequality. We use the notation $\mathbb{P}\otimes\phi_{p,\Lambda_{N}}^{\gamma,0}$ to denote the product measure of the external field and the FK-Ising measure, where we emphasize that the second component is simply the FK-Ising measure without disorder. Let $\mathcal{A}_{2},\mathcal{A}_{3}\subset\mathbb{R}^{\Lambda_{N}}\otimes\{0,1\}^{\rm{E}(\Lambda_{N})}$ denote the events that
$\sum_{\mathcal{C}\in\mathfrak{C}}f\left({\frac{\epsilon h_{\mathcal{C}}}{T}}\right)-\frac{\epsilon^{2}N^{2}}{2T^{2}}\leq\epsilon N^{\alpha(T)}~{}~{}\mbox{and}~{}~{}\sum_{\mathcal{C}\in\mathfrak{C}}f\left({\frac{\epsilon h_{\mathcal{C}}}{T}}\right)-\frac{\epsilon^{2}N^{2}}{2T^{2}}\geq-\epsilon N^{\alpha(T)}.$
Then we have (for $i=2,3$):
$\mathbb{P}\otimes\phi_{p,\Lambda_{N}}^{\gamma,0}(\mathcal{A}_{i}^{c})\,\geq\,\mathbb{P}\otimes\phi_{p,\Lambda_{N}}^{\gamma,0}\left(\mathcal{A}_{i}^{c}|(\mathcal{H}^{i})^{c}\right)\times\mathbb{P}\left((\mathcal{H}^{i})^{c}\right)\geq\exp\left(-\epsilon^{-\frac{1}{4}}N^{-\frac{\alpha(T)}{4}}\right)\times\mathbb{P}\left((\mathcal{H}^{i})^{c}\right)\,.$ (4.15)
Therefore, all we need is to show that $\mathcal{A}_{i}^{c}$ has a small probability under $\mathbb{P}\otimes\phi_{p,\Lambda_{N}}^{\gamma,0}$.
For any $\omega$ fixed, we write $h_{\mathcal{C}}=|\mathcal{C}|X_{\mathcal{C}}$, then we get that $X_{\mathcal{C}}~{}(\mathcal{C}\in\mathfrak{C})$’s are i.i.d. Gaussian variables with mean 0 and variance 1 and denote the law as $\mathbb{P}_{X}$. Now we can rewrite the second moment term in (4.12) as
$\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}=\sum_{\mathcal{C}\in\mathfrak{C}(\omega)}\frac{\epsilon^{2}|\mathcal{C}|}{2T^{2}}(X_{\mathcal{C}}^{2}-1)+\frac{\epsilon^{2}N^{2}}{2T^{2}}.$
The first term $\sum_{\mathcal{C}\in\mathfrak{C}(\omega)}\frac{\epsilon^{2}|\mathcal{C}|}{2T^{2}}(X_{\mathcal{C}}^{2}-1)$ is a centralized summation of $\chi$-square distributions which can be estimated using [27, Lemma 1]. Let $\mathscr{S}=\{\omega:\sum_{\mathcal{C}\in\mathfrak{C}{(\omega)}}|\mathcal{C}|^{2}\leq\epsilon^{-1}N^{3\alpha(T)}\}.$ Then for any $\omega\in\mathscr{S}$, we have by (4.12)
@@
The last inequality follows from the exponential tail bound for $\chi$-square distributions ([27, Lemma 1]). Similarly, we have by (4.12)
@@
Again, by the exponential tail bound for $\chi$-square distributions ([27, Lemma 1]), we have
${\mathbb{P}_{X}}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}|\mathcal{C}|}{2T^{2}}(X_{\mathcal{C}}^{2}-1)<-\frac{1}{2}\epsilon N^{\alpha(T)}\right)\leq C_{1}\exp\left(-C_{1}^{-1}\epsilon^{-\frac{1}{2}}N^{-\frac{\alpha(T)}{2}}\right).$ (4.18)
In addition, by Lemma B.5, we have
${\mathbb{P}_{X}}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{4}|\mathcal{C}|^{2}}{2T^{4}}X_{\mathcal{C}}^{4}>\frac{1}{2}\epsilon N^{\alpha(T)}\right)\leq C_{1}\exp\left(-C_{1}^{-1}\epsilon^{-\frac{1}{2}}N^{-\frac{\alpha(T)}{2}}\right).$ (4.19)
@@
for $i=2,3,$ some $C_{3}>0$ and sufficiently large $N$. Combining (4.22) with (4.15), we complete the proof of the lemma.
@@
Let $\mathcal{H}=\mathcal{H}^{0}\cap\mathcal{H}^{1}\cap\mathcal{H}^{2}\cap\mathcal{H}^{3}$, then combining Lemmas 4.3, 4.6 and 4.8 shows
$\mathbb{P}(\mathcal{H})\geq 1-C_{1}\exp\left(-C_{1}^{-1}\epsilon^{-\frac{1}{2}}N^{-\frac{\alpha(T)}{2}}\right)$
for some constant $C_{1}>0$. For $h\in\mathcal{H}$, by (4.13) and (4.14), we compute that
$\phi_{p,\Lambda_{N}}^{\gamma,0}\left(\exp(-\epsilon N^{\alpha(T)})\leq\frac{\prod_{\mathcal{C}\in\mathfrak{C}}\cosh(\frac{\epsilon h_{\mathcal{C}}}{T})}{\exp(\frac{2\epsilon^{2}N^{2}}{T^{2}})}\leq\exp\left(\epsilon N^{\alpha(T)}\right)\right)\geq 1-2\exp\left(-\epsilon^{-\frac{1}{4}}N^{-\frac{\alpha(T)}{4}}\right).$ (4.23)

### v2

@@ 4 Absolute continuity in weak-disorder regime
is close to 1 for typical disorder when $\epsilon\ll N^{-\beta(T)}$ where
@@
In Section 4.1, we show the concentration of $Z(h)$ using a similar technique as in [12, Theorem 3.14] and in Section 4.2, we apply Lemma 3.1 to show the absolute continuity.
@@ 4.1 Concentration of $Z(h)$
Let $\Omega$ be a region such that $\Lambda_{N}\subset\Omega\subset\Lambda_{2N}$ and $\gamma\in\{-1,1\}^{\partial_{int}\Omega}$ be a boundary condition on $\partial_{int}\Omega$. Let $\mathcal{A}_{0}\subset{\mathbb{R}}^{\Omega}\otimes\Xi$ denote the collection of pairs $(h,\omega)$ such that the property (P0) holds,
$(\text{\emph{P}}0):\quad\Bigg{|}\Bigg{\langle}~{}\prod_{v\in\Omega}\Bigg{[}1+\sigma_{v}\tanh\left(\frac{\epsilon h_{v}}{T}\right)\Bigg{]}~{}\Bigg{\rangle}^{\gamma,0}_{T,\Omega}-1\Bigg{|}\leq\sqrt{\epsilon N^{\beta(T)}}.$ (4.4)
Note that actually $\mathcal{A}_{0}$ depends only on $h$, so we could define $\mathcal{A}_{0}\subset{\mathbb{R}}^{\Omega}$ to be the collection of the external field $h$ such that the property (P0) holds. However, for further convenience, we keep $\omega$ in the definition of $\mathcal{A}_{0}$.
@@
Of course, the set $\mathcal{A}_{0}$ relies on $\gamma,\Omega,T$ and $\epsilon,$ but we omit them from the notation for concision. The same conventions will be used several times without further notice.
@@
Take $\Omega=\Lambda_{N},$ then there exist constants $c_{1},c_{2}>0$ depending only on $T$ such that $\mathbb{P}\otimes\phi_{p,\Omega}^{\gamma,0}(\mathcal{A}^{0})\geq 1-c_{1}\exp\left(-c_{1}^{-1}\sqrt{\epsilon^{-1}N^{-\beta(T)}}\right)$ for any disorder strength $\epsilon\leq c_{2}N^{-\beta(T)}$ and any boundary condition $\gamma$ on $\partial_{int}\Lambda_{N}$.
@@
The case $T=T_{c}$ is covered by [15, Lemma 3.4]. The cases that $T<T_{c}$ and $T>T_{c}$ follow a similar proof strategy. Here, we only highlight the key adaptations required. Note that
@@
To control the right-hand side, using [15, Lemma B.1], the core is to bound
@@
For $T>T_{c}$, by the Edward-Sokal coupling between the Ising model and the FK-Ising model, we have $\langle\sigma^{J}\rangle^{\gamma,0}_{T,\Lambda_{N}}=\phi^{\gamma,0}_{p,\Lambda_{N}}(\mathtt{Even}_{J}),$ where $\mathtt{Even}_{J}$ stands for the event that each cluster in $\omega^{\gamma}$ contains an even number of points in $J$.
@@
For any $J\subset\Lambda_{N}$, let $d_{x}=\frac{1}{2}\cdot dist(x,\partial_{int}\Lambda_{N}\cup J\setminus\{x\})$ and let $\mathcal{R}_{x}=\Lambda_{d_{x}}(x)$. Then it is easy to get that
$\mathtt{Even}_{J}\subset\bigcap_{x\in J}\{x\longleftrightarrow\partial_{int}\mathcal{R}_{x}\}.$ (4.6)
@@
$\langle\sigma^{J}\rangle^{\gamma,0}_{T,\Lambda_{N}}\leq\prod_{x\in J}\phi_{p,\mathcal{R}_{x}}^{\mathrm{w},0}\left(\{x\longleftrightarrow\partial_{int}\mathcal{R}_{x}\}\right)\leq\exp{\Big{(}-\sum_{x\in J}c\cdot d_{x}\Big{)}}.$ (4.7)
Then we get from the calculation in [12, Lemmas 8.1 and 8.3] that
@@
With the upper bound for $A_{k}$, the rest of the proof is the same to that in [15, Lemma 3.4].
@@
If $T\geq T_{c}$ and $\gamma=\mathrm{f}$ is the free boundary, then we get from CBC that $\langle\sigma^{J}\rangle^{\mathrm{f},0}_{T,\Omega}\leq\langle\sigma^{J}\rangle^{\mathrm{w},0}_{T,\Lambda_{2N}}.$ The desired bound on $A_{k}$ then follows from the proof of Lemma 4.3. ∎
@@
Let $\mathcal{A}_{1}\subset{\mathbb{R}}^{\Omega}\otimes\Xi$ denote the collection of pairs $(h,\omega)$ such that the property (P1) holds
@@
There exist constants $c_{1}>0$ and $c_{2}>0$ depending only on $T$ such that for any disorder strength $\epsilon\leq c_{1}N^{-\frac{1}{2}}$, we have $\mathbb{P}\otimes\phi_{p,\Omega}^{\gamma,0}(\mathcal{A}_{1})\geq 1-c_{2}\exp\left(-c_{2}^{-1}\epsilon^{-1}N^{-\frac{1}{2}}\right)$.
@@
Recall that $f(x)=\ln\cosh(x)$. Given the inequality $\frac{x^{2}-x^{4}}{2}\leq f(x)\leq\frac{x^{2}}{2}$, we calculate that
@@
Combining (4.3), (4.4) and (4.9), we obtain that for $(h,\omega)\in\mathcal{A}_{0}\cap\mathcal{A}_{1},$
$\left(1-\sqrt{\epsilon N^{\beta(T)}}\right)\left(1-\sqrt{\epsilon^{2}N}\right)$ $\leq Z(h)\cdot\exp\left(\frac{2\epsilon^{2}N^{2}}{T^{2}}\right)$ $\leq\left(1+\sqrt{\epsilon N^{\beta(T)}}\right)\left(1+\sqrt{\epsilon^{2}N}\right).$ (4.11)
@@ 4.2 Concentration of $F(h,\omega)$
Combined with Lemma 3.1, we can consider the “typical” event such that $F(h,\omega)$ concentrates well under the measure $\mathbb{P}\otimes\phi_{p,\Lambda_{N}}^{\gamma,0}$.
@@
We denote by $\mathcal{A}_{2}$ and $\mathcal{A}_{3}$ the collections of pairs $(h,\omega)$ such that the following properties (P2) and (P3) hold respectively.
$(\text{\emph{P}}2):\quad F(h,\omega)-\frac{2\epsilon^{2}N^{2}}{T^{2}}\leq\epsilon N^{\alpha(T)},$ (4.13)
$(\text{\emph{P}}3):\quad F(h,\omega)-\frac{2\epsilon^{2}N^{2}}{T^{2}}\geq-\epsilon N^{\alpha(T)}.$ (4.14)
@@
$\mathbb{P}\otimes\phi_{p,\Lambda_{N}}^{\gamma,0}(\mathcal{A}_{2}\cap\mathcal{A}_{3})\geq 1-c_{2}\exp\big{(}-c_{2}^{-1}\epsilon^{-\frac{1}{2}}N^{-\frac{\alpha(T)}{2}}\big{)}.$
@@
For any $\omega$ fixed, we write $X_{\mathcal{C}}=\frac{h_{\mathcal{C}}}{\sqrt{|\mathcal{C}|}}$, then we get that $X_{\mathcal{C}}~{}(\mathcal{C}\in\mathfrak{C})$’s are i.i.d. Gaussian variables with mean 0 and variance 1 and we denote the law as $\mathbb{P}_{X}$. Now we can rewrite the second moment term in (4.12) as
$\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}h_{\mathcal{C}}^{2}}{2T^{2}}=\sum_{\mathcal{C}\in\mathfrak{C}(\omega)}\frac{\epsilon^{2}|\mathcal{C}|}{2T^{2}}(X_{\mathcal{C}}^{2}-1)+\frac{2\epsilon^{2}N^{2}}{T^{2}}.$
The first term $\sum_{\mathcal{C}\in\mathfrak{C}(\omega)}\frac{\epsilon^{2}|\mathcal{C}|}{2T^{2}}(X_{\mathcal{C}}^{2}-1)$ is a centralized summation of $\chi$-square distributions which can be estimated using Lemma B.5. Let $\mathscr{S}=\{\omega:\sum_{\mathcal{C}\in\mathfrak{C}{(\omega)}}|\mathcal{C}|^{2}\leq\epsilon^{-1}N^{3\alpha(T)}\}.$ Then for any $\omega\in\mathscr{S}$, we have by (4.12)
@@
The last inequality follows from Lemma B.5. Similarly, we have by (4.12)
@@
Again by Lemma B.5, we have
${\mathbb{P}_{X}}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{2}|\mathcal{C}|}{2T^{2}}(X_{\mathcal{C}}^{2}-1)<-\frac{1}{2}\epsilon N^{\alpha(T)}\right)$ $\leq C_{1}\exp\left(-C_{1}^{-1}\epsilon^{-\frac{1}{2}}N^{-\frac{\alpha(T)}{2}}\right),$ (4.17)
and $\mbox{and }~{}~{}~{}{\mathbb{P}_{X}}\left(\sum_{\mathcal{C}\in\mathfrak{C}}\frac{\epsilon^{4}|\mathcal{C}|^{2}}{2T^{4}}X_{\mathcal{C}}^{4}>\frac{1}{2}\epsilon N^{\alpha(T)}\right)$ $\leq C_{1}\exp\left(-C_{1}^{-1}\epsilon^{-\frac{1}{2}}N^{-\frac{\alpha(T)}{2}}\right).$ (4.18)
@@
for $i=2,3,$ some $C_{3}>0$ and sufficiently large $N$. Thus we complete the proof. ∎
@@
Let $\mathcal{A}=\mathcal{A}_{0}\cap\mathcal{A}_{1}\cap\mathcal{A}_{2}\cap\mathcal{A}_{3}$, then combining Lemmas 4.3, 4.6 and 4.8 shows
$\mathbb{P}\otimes\phi_{p,\Lambda_{N}}^{\gamma,0}(\mathcal{A})\geq 1-C_{1}\exp\left(-C_{1}^{-1}\epsilon^{-\frac{1}{2}}N^{-\frac{\alpha(T)}{2}}\right)$
for some constant $C_{1}>0$. For $(h,\omega)\in\mathcal{A}$, by (4.13) and (4.14), we compute that
$\exp(-\epsilon N^{\alpha(T)})\leq\frac{\prod_{\mathcal{C}\in\mathfrak{C}}\cosh(\frac{\epsilon h_{\mathcal{C}}}{T})}{\exp(\frac{2\epsilon^{2}N^{2}}{T^{2}})}\leq\exp\left(\epsilon N^{\alpha(T)}\right).$ (4.22)